ERAM 30 x-ray scan is done.

RH at 0,2 % after 15 h at 16l/h.

Mounting of Argon supply circuit this week.

mechanics is not seen, but need to investigate high gain region and the mean gain is low (~1200)

6 bottles received. 6 new to be ordered this week

order to be done this week

Full scan this week semaine (25.07):

24 is planned this week

_____________________________________________

29 in priority

26

20

the last 400 kOhm one

11.

redo 23 because we have two half scans

redo 13 because we are not sure we are using the right cards

redo 17